# Network-Realised Model Predictive Control Part II: Distributed Constraint Management

### v4

@@ 6.2 Guaranteeing Recursive Feasibility
In order to extend the theoretical guarantees obtained in Proposition 6.1 beyond the initial time instant, we must ensure that each of the MPC-based control laws from (8) is feasible for every time instant $k\in\mathbb{N}$ with $k\geq k_{0}$. This property, known as recursive feasibility (see Section 12.3 in [BBM]), plays a key role in our chosen approach and, consequently, the following theorem (which directly builds upon the results obtained in Proposition 6.1) represents the paper’s main theoretical result.
@@
For each of the area dynamics described in (16a)-(16d), let an MPC subcontroller be implemented as in Algorithm 1 and let the communicated information between these subcontrollers be affected as in (20). For all $i\in\{1:N\}$, let $\rho_{i}\in\mathbb{N}$ with $\rho_{i}\geq 1$ and define
@@
$\qquad\quad\left.p_{uj}\in{\mathcal{P}}_{2(j+\alpha_{ui})t},\,j\in\{1:n_{ui}\}\right\}.$ (45b)
Assume that:
@@ 2. A2)
    For each $i\in\{1:N\}$, there exists $\rho_{i}\in\mathbb{N}$ with $\rho_{i}\geq 1$ which satisfies
@@
Notice the fact that it is possible to retrieve the above-mentioned guarantees by taking $\rho_{i}=1$, for every $i\in\{1:N\}$, resulting in especially efficient implementations (see the numerical example from the sequel) from a computational standpoint. Indeed, as the values of $\rho_{i}$ increase, not only does (46) become more difficult to satisfy, but the additional constraints associated with taking $T_{i}=\rho_{i}$ may dramatically increase the cost of solving (8), depending on the complexity of the sets in (44). As discussed in Remark 6.4, taking larger values for $\overline{T}_{i}$ may yield similar degrees of performance as increasing $T_{i}$, without greatly impacting overall computational costs.
To summarise the design workflow of our architecture’s second layer, we provide a high-level algorithm (located at the top of the next page), which condenses the various topics discussed in this paper into a sequence of steps.
@@
Performing Steps 3a-3c in Algorithm 2 makes the paramount importance of our architecture’s first layer readily apparent. It is precisely due to the proper tuning of the NRF-based control laws that area cross-coupling may be diminished (and potentially rendered negligible, for all practical purposes) to such a degree that the MPC-based subcontrollers may be tuned individually via (30a)-(30b). While this effectively bypasses the need to carefully construct the sets from (30a)-(30b) and it crucially prevents the design procedure of the MPC-based subcontrollers from being co-dependent in any significant way, we point out that said procedure does indeed require information pertaining to the first-layer response of other areas. However, this information is fully available once the NRF-based control laws have been obtained, and it is precisely the role of these distributed controllers to ensure the feasibility (recall Remark 6.2) and the relaxed co-dependency of the second layer’s design.
@@
Step 4a: Try increasingly larger values for $\rho_{i}\geq 2$ and then check (46) for all $i\in\{1:N\}$;
@@ 6.3 Computational Considerations
The following discussion is centred around the three core concepts which underpin the computational burden of our procedure, namely: synthesis costs, design scalability, and implementation complexity.
@@ 6.3.1 Synthesis Costs
One of the main features associated with Algorithm 2 is the fact that all of its constituent steps (along with the design phase of the first layer, see our companion paper [part1]) can be performed offline, and that any reconfiguration of the second layer’s constraints or cost functions does not impact the first layer’s control laws (in contrast to the technique from [DMPC1, DMPC2], recalling Remark 6.1). The chief computational burden associated with our solution is represented by the cost of computing the sets from (34a)-(34b) along with (42)-(43) for large-scale applications. Although the measures described in Remarks 5.5 and 6.3 can alleviate some of the complexity associated with these operations, the latter remain heavily dependent on the shape and the structure of the involved sets, making it challenging to come up with a one-size-fits-all computational cost. In spite of this, switching to a homothetic [homot] or even zonotope-based [zonot] framework (while employing various approximations, either in terms of the operation or of the operands themselves) is guaranteed to have a significant, positive impact on the complexity of the synthesis procedure. For a practical and specific example of the costs associated with the design of the second layer, see the numerical example presented in the sequel, and the source files from its GitHub repository.
@@ 6.3.2 Design Scalability
While all of the steps which form Algorithm 2 can, indeed, be iterated in a fully offline setting, yet another appealing feature is the fact that once the second layer’s command budgeting has been negotiated via the selection of the sets from (30a)-(30b), the subsequent set-based computations from Algorithm 2 can be performed in parallel. As indicated in Remark 6.7, the design procedure of each area can proceed independently of the others, once the aforementioned choice has been made. This fact is a direct consequence of not employing any notions of joint set invariance (see, for example, [R8C]) and of relying, instead, upon the decoupling-focused control laws discussed in our companion paper [part1].
@@ 6.3.3 Implementation Complexity
Due to the separate implementations in (10) and in Algorithm 1, a crucial observation is the fact that the runtime cost of our architecture is divided among the network’s areas. Said burden is appealingly modest, since it reduces to several inexpensive operations, namely: four matrix-vector multiplications and two vector-vector additions for an area’s first-layer subcontroller (10), along with the solving of (8) for the second layer. We point out that the latter problem boils down to a standard Quadratic Programming (QP) optimisation problem, whenever the cost in (8) has a quadratic (and symmetrical) expression and the sets from (30a)-(30b) are chosen to be polyhedral in nature (which is always possible, as per Remark 5.3). Despite the fact that the overall complexity of this problem is highly dependent upon the number of constraints it includes (recall Remark 6.3), note that numerous solvers and optimisation toolboxes (such as [mosek] and [YALMIP], respectively) are widely available for the efficient solving of such problems.
@@ 7.1.1 The Dynamics of the Application
In order to showcase the effectiveness of the proposed design framework, we now focus on a numerical example centred around the vehicle platoon application discussed in [plutonizare]. We consider a formation of $N=10$ identical vehicles, in which the continuous-time models from (22) in [plutonizare] are configured with the following parameters: $m_{i}=1$, $\tau_{i}=0.1$ and $\sigma_{i}=1,\,\forall\,i\in\{1:N\}$. The resulting zero-order hold discretised dynamics of each car, for a sampling period of $T_{s}=0.1$, are given by
@@
where $u_{i}$ is the control signal of the $i^{\text{th}}$ car, $p_{i}[k]$ denotes its (absolute) position, $v_{i}[k]$ is its speed and $\mu_{i}[k]$ is the state of the $i^{\text{th}}$ car’s actuator (the $\Phi$ terms in [plutonizare]).
@@ 7.1.2 First-Layer Objectives and Platoon Model
In keeping with the control laws proposed in [plutonizare], we wish to designate each car as an individual area and to obtain a first layer in which the control signals are computed (in the absence of exogenous or second-layer signals) as
@@
for all $i\in\{2:N\}$, where $y_{i}[k]:=p_{i}[k]+\ell_{i}[k]-p_{i-1}[k]$ stands for the $i^{\text{th}}$ car’s length-based interdistance and $h>0$ is the platoon’s time headway (see [plutonizare] and the references therein for its theoretical justification), which has been chosen as $h=5$ for the considered application.
@@
$T=\tiny\begin{bmatrix}I_{N+1}&O&O&O&\dots&O&O\\ T_{\ell 1}&I_{3}&O&O&\dots&O&O\\ T_{\ell 2}&T_{\ell d}&I_{3}&O&\dots&O&O\\ T_{\ell 3}&O&T_{\ell d}&I_{3}&\dots&O&O\\ \vdots&\vdots&&\ddots&\ddots&\vdots&\vdots\\ T_{\ell N}&O&O&O&\dots&T_{\ell d}&I_{3}\end{bmatrix}\in\mathbb{R}^{(4N+1)\times(4N+1)},\vskip-8.53581pt$
@@
By choosing $\mathbb{C}_{g}=\{z\in\mathbb{C}\ |\ |z|<1\}$, the first layer will aim to stabilise platoon’s closed-loop dynamics, in the sense that, by implementing the control laws of type (47) as detailed in our companion paper [part1], the closed-loop dynamics will be described by
@@ 7.1.3 Second-Layer Objectives
These constraints must be satisfied while employing the same communication scheme as the one considered in (47), i.e., the $i^{\text{th}}$ car can employ only information that is accessed locally or from the car directly in front of it.
Additionally, due to the fact that the prototype NRF controller from [plutonizare] (if properly tuned) ensures a high level of performance, we want the second layer to intervene as little as possible in the closed-loop dynamics of the first one. In other words, most of the control action should be undertaken by the $u_{fi}[k]$ signals, whereas the entries of $u_{s1i}[k]$ and $u_{s2i}[k]$ should have negligible absolute values (compared to those of $u_{fi}[k]$) whenever the constraints from (50) are not in danger of being violated.
@@ 7.2.1 The First Layer
Since the application considered in this numerical example already has a dedicated NRF-based solution, we defer to [plutonizare] for the procedural details regarding the design of our first layer. In so doing, we briefly discuss the numerical implementation of these control laws, before focusing our attention on the design of the second layer.
@@
By using the state-space systems in (48) and (51)-(52) to form the first layer’s closed-loop system, we obtain a realisation of type (49) in which the spectral radius of $A_{f}$ is $0.9936$, thus validating all of the design objectives associated with the first layer. In addition to this, the resulting closed-loop dynamics of each car in the platoon will be given by the realisations from (53a)-(53b), which are located at the top of this page and in which
$A_{w}=\left[\scriptsize\begin{array}[]{ccc}0&\ -0.1&\ 0.0331\\ 0&\ 0&\ 0\\ 0&\ 0&\ 0\\ \end{array}\right]\text{ and }B_{w}=\left[\scriptsize\begin{array}[]{c}-0.0381\\ 0\\ 0\\ \end{array}\right].\vskip-8.53581pt$
@@ 7.2.2 The Second Layer
The first step in designing our control architecture’s second layer, with the aim of ensuring its recursive feasibility, consists of defining all of the sets discussed in Section 5. We begin by addressing the command signals
generated by the MPC layer, for which we impose:
@@
and, finally, we address the constraints related to the platoon’s states by first introducing the matrix denoted $W_{x}:=\mathrm{row}_{1}\left(\tiny\begin{bmatrix}A_{car}&B_{car}\end{bmatrix}\right)\mathrm{diag}(0,I_{3})$ and then constructing the set $\widehat{\mathcal{X}}\subseteq\mathbb{R}^{4}$ given explicitly in (54), which is located at the top of this page.
@@
As explained in the comments that accompany the source code from the above-indicated link, our approach entails the implicit design of the robust control invariant sets $\mathcal{C}_{i}\subseteq\widehat{\mathcal{X}}$, which (by their very construction; see, for example, Section 10.4 in [BBM]) guarantee the recursive feasibility of our second layer. We stress the fact that (as previously indicated in Section 6) the sets $\mathcal{C}_{i}$ are not computed explicitly, with their existence being a desirable and natural consequence of our proposed approach.
@@ 7.3 Simulation Scenario and Results
In order to test the obtained control laws, we simulate a scenario in which $\Delta p_{0}[k]=T_{s}v_{0}[k]$, where
@@
is the speed of the virtual $0^{\text{th}}$ car and $\mathds{1}[k]$ denotes the discrete-time Heaviside step function. As can be seen in Figure 3 below, in which $v_{0}[k]$ is depicted in orange, the speed of each car in the platoon will try to match that of the $0^{th}$ car in order to achieve steady-state evolution.
@@
Moreover, as can be seen in Figure 4, our two-layer control architecture successfully maintains all of the interdistance variables within their prescribed bounds, thus avoiding collisions and preventing the cars from drifting too far apart (leading to traffic congestion or the breakdown of the communication between consecutive cars).
@@
With respect to the limitations placed upon the signals sent to the car actuators, our control scheme is once again successful. As can be seen from Figures 5 and 6 below, the sum of the signals from the two aforementioned figures is well within the bounds from (50), even taking into account the amplitudes of the entries in $\beta_{s2}[k]$.
@@
The brief intervals in which the two signal vectors are of the same order of magnitude as the first-layer control signals are those in which the second layer intervenes in order to saturate either the output of the first car’s NRF-based subcontroller (see Figure 8 below), or the speed of the lead car in the platoon (see Figure 9, also located below) with respect to these variables’ prescribed bounds.
@@
Finally, as anticipated in Section 6.3.3 and as showcased by the runtime statistics showcased in Tables 3 and 4 above, the implementation costs associated with our proposed solution are sufficiently modest such that command computation is well below the sampling period of our control system, namely $T_{s}=0.1$ seconds. Note that all simulations associated with these runtimes were performed via MATLAB R2024b, on a mid-range laptop equipped with an i9-100885H Intel Core processor.

### v5

@@ 6.2 Guaranteeing Recursive Feasibility
In order to extend the theoretical guarantees obtained in Proposition 6.1 beyond the initial time instant, we must ensure that each of the MPC-based control laws from (8) is feasible for every time instant $k\in\mathbb{N}$ with $k\geq k_{0}$. This property, known as recursive feasibility (see Section 12.3 in [5]), plays a key role in our chosen approach and, consequently, the following theorem (which directly builds upon the results obtained in Proposition 6.1) represents the paper’s main theoretical result.
@@
For each of the area dynamics described in (16a)-(16d), let an MPC subcontroller be implemented as in Algorithm 1 and let the communicated information between these subcontrollers be affected as in (20). For all $i\in\{1:N\}$, define also the following sets:
@@
$\qquad\quad\left.p_{uj}\in{\mathcal{P}}_{2(j+\alpha_{ui})t},\,j\in\{1:n_{ui}\}\right\},$ (45b)
and assume that:
@@ 2. A2)
    For all $i\in\{1:N\}$, there exists $\rho_{i}\in\mathbb{N}$ with $\rho_{i}\geq 1$, such that the following inclusions hold:
@@
Notice the fact that it is possible to retrieve the above-mentioned guarantees by taking $\rho_{i}=1$, for every $i\in\{1:N\}$, resulting in exceptionally efficient implementations (see the numerical example from the sequel) from a computational standpoint. Indeed, as the values of $\rho_{i}$ increase, not only does (46) become more challenging to satisfy, but the additional constraints associated with taking $T_{i}=\rho_{i}$ may dramatically increase the cost of solving (8), depending on the complexity of the sets in (44). As discussed in Remark 6.4, taking larger values for $\overline{T}_{i}$ may yield similar degrees of performance as increasing $T_{i}$, without significantly impacting computational costs.
To summarise the design workflow of our architecture’s second layer, we provide a high-level algorithm (located at the top of the next page) that condenses the myriad topics discussed in this paper into a sequence of steps.
@@
Performing Steps 3a-3d in Algorithm 2 makes the paramount importance of our architecture’s first layer readily apparent. It is precisely due to the proper tuning of the NRF-based control laws that area cross-coupling may be diminished (and potentially rendered negligible, for all practical purposes) to such a degree that the MPC-based subcontrollers may be tuned individually via (31a)-(31b). While this fact effectively bypasses the need to carefully construct the sets from (31a)-(31b) and it crucially prevents the design procedure of the MPC-based subcontrollers from being co-dependent in any significant way, we point out that said procedure does indeed require information related to the first-layer response of other areas. However, this information becomes fully available once the NRF-based control laws have been obtained, and it is precisely the role of these distributed controllers to ensure the feasibility (recall Remark 6.2) and relaxed co-dependency of the second layer’s design.
@@
Step 4a: For all $i\in\{1:N\}$, try increasingly larger values for $\rho_{i}\geq 2$ and then check (46);
@@ 6.3 Computational Considerations
The following discussion revolves around the three core concepts that underpin the computational burden of our procedure: synthesis costs, design scalability, and implementation complexity.
@@ 6.3.1 Synthesis Costs
One of the main features associated with Algorithm 2 is the fact that all of its constituent steps (along with the design phase of the first layer, see our companion paper [24]) can be performed offline, and that any reconfiguration of the second layer’s constraints or cost functions does not impact the first layer’s control laws (in contrast to the technique from [1, 2], recalling Remark 6.1 ). The chief computational burden associated with our solution is the cost of computing the sets in (34a)-(34b), along with those in (42)-(43), for large-scale applications. Although the measures described in Remarks 5.5 and 6.3 can alleviate some of the complexity associated with these operations, the latter remain heavily dependent on the shape and the structure of the involved sets, making it challenging to come up with a one-size-fits-all computational cost. In spite of this fact, switching to a homothetic [18] or even zonotope-based [13] framework (while employing various approximations, either in terms of the operation or of the operands themselves) is guaranteed to have a significant, positive impact on the complexity of the synthesis procedure. For a practical and specific example of the costs associated with the design of the second layer, see the numerical example presented in the sequel, and the source files from its GitHub repository.
@@ 6.3.2 Design Scalability
While all of the steps which form Algorithm 2 can, indeed, be iterated in a fully offline setting, yet another appealing feature is the fact that once the second layer’s command budgeting has been negotiated via the selection of the sets from (31a)-(31b), the subsequent set-based computations from Algorithm 2 can be performed in parallel. Thus, the design procedure of each area can proceed independently of the others, once the aforementioned choice has been made. This fact is a direct consequence of not employing any notions of joint set invariance (see, for example, [22]) and of relying, instead, upon the decoupling-focused control laws discussed in [24].
@@ 6.3.3 Implementation Complexity
Due to the separate implementations described in (10) and in Algorithm 1, a key feature of our proposed solution is the fact that the runtime cost of our architecture is divided among the network’s areas. Moreover, said computational burden is appealingly modest, since it reduces to several inexpensive operations, namely: four matrix-vector multiplications and two vector-vector additions for an area’s first-layer subcontroller (10), along with the solving of the optimisation problem from (8) for each area’s second-layer subcontroller. In addition to this, we point out that each of the latter problems boils down to a standard Quadratic Programming (QP) optimisation problem, whenever the cost in (8) has a quadratic (and symmetrical) expression and the sets from (31a)-(31b) are chosen to be polyhedral in nature (which is always possible, as per Remark 5.3). Although the overall complexity of this problem is highly dependent upon the number of constraints it includes (recall Remark 6.3), note that numerous solvers and optimisation-based toolboxes (such as [3] and [14], respectively) are widely available for the efficient solving of such standard problems.
@@ 7.1.1 The Dynamics of the Application
To showcase the effectiveness of the proposed design framework, we now focus on a numerical example centred on the vehicle platoon application discussed in [20]. We consider a formation of $N=10$ identical vehicles, in which the continuous-time models from (22) in [20] are configured with the following parameters: $m_{i}=1$, $\tau_{i}=0.1$ and $\sigma_{i}=1,\,\forall\,i\in\{1:N\}$. The resulting zero-order-hold-discretised dynamics of each car, for a sampling period of $T_{s}=0.1$, are given by
@@
where $u_{i}[k]$ is the control signal of the $i^{\text{th}}$ car, $p_{i}[k]$ denotes its (absolute) position, $v_{i}[k]$ is its speed and $\mu_{i}[k]$ is the state of the $i^{\text{th}}$ car’s actuator (the $\Phi$ terms in [20]).
@@ 7.1.2 First-Layer Objectives and Platoon Model
In keeping with the control laws proposed in [20], we wish to designate each car as an individual area and to obtain a first layer in which the control signals are computed (in the absence of exogenous or second-layer signals) as
@@
for all $i\in\{2:N\}$, where $y_{i}[k]:=p_{i}[k]+\ell_{i}[k]-p_{i-1}[k]$ stands for the $i^{\text{th}}$ car’s length-based interdistance and $h>0$ is the platoon’s time headway (see [20] and the references therein for its theoretical justification), which has been chosen as $h=5$ for the considered application.
@@
$T=\scriptsize\begin{bmatrix}I_{N+1}&O&O&O&\dots&O&O\\ T_{\ell 1}&I_{3}&O&O&\dots&O&O\\ T_{\ell 2}&T_{\ell d}&I_{3}&O&\dots&O&O\\ T_{\ell 3}&O&T_{\ell d}&I_{3}&\dots&O&O\\ \vdots&\vdots&&\ddots&\ddots&\vdots&\vdots\\ T_{\ell N}&O&O&O&\dots&T_{\ell d}&I_{3}\end{bmatrix}\in\mathbb{R}^{(4N+1)\times(4N+1)},\vskip-8.53581pt$
@@
By choosing $\mathbb{C}_{g}=\{z\in\mathbb{C}\ |\ |z|<1\}$, the first layer will aim to stabilise platoon’s closed-loop dynamics, in the sense that, by implementing the control laws of type (47) as detailed in our companion paper [24], the closed-loop
dynamics will be described by
@@ 7.1.3 Second-Layer Objectives
These constraints must be satisfied while employing the same communication scheme as in (47), i.e., the $i^{\text{th}}$ car can use only information available locally (from its own readings) or from the car directly in front of it.
Additionally, due to the fact that the prototype NRF-based controllers discussed in [20] (when properly tuned) achieve high performance, we want the second layer to intervene as little as possible in the closed-loop dynamics of the first. In other words, most of the control action should be undertaken by the $u_{fi}[k]$ signals, whereas the entries of $u_{s1i}[k]$ and $u_{s2i}[k]$ should have negligible absolute values (compared to those of $u_{fi}[k]$) whenever the constraints from (50) are not in danger of being violated.
@@ 7.2.1 The First Layer
Since the application considered in this numerical example already has a dedicated NRF-based solution, we defer to [20] for the procedural details of our first-layer design. In so doing, we briefly discuss the numerical implementation of these control laws before focusing our attention on the design of the second layer.
@@
By using the state-space systems in (48) and (51)-(52) to form the first layer’s closed-loop system, we obtain a realisation of type (49) in which the spectral radius of $A_{f}$ is $0.9936$, thus validating all of the design objectives associated with the first layer. In addition to this fact, the resulting closed-loop dynamics of each car in the vehicle platoon will be given by the realisations stated in (53a)-(53b), located at the top of this page and where
$A_{w}=\left[\begin{array}[]{ccc}0&\ -0.1&\ 0.0331\\ 0&\ 0&\ 0\\ 0&\ 0&\ 0\\ \end{array}\right]\text{ and }B_{w}=\left[\begin{array}[]{c}-0.0381\\ 0\\ 0\\ \end{array}\right].\vskip-8.53581pt$
@@ 7.2.2 The Second Layer
The first step in designing our second-layer control system, with the aim of ensuring its recursive feasibility, is to define all the sets discussed in Section 5. We begin by addressing the command signals generated by the MPC layer, for which we impose the following constraints:
@@
and, finally, we address the constraints related to the platoon’s states by first introducing the matrix denoted $W_{x}:=\mathrm{row}_{1}\left(\tiny\begin{bmatrix}A_{car}&B_{car}\end{bmatrix}\right)\mathrm{diag}(0,I_{3})$ and then constructing the set $\widehat{\mathcal{X}}\subseteq\mathbb{R}^{4}$ given in (54) at the top of this page.
@@
As explained in the comments that accompany the source code from the above-indicated link, our approach entails the implicit design of the robust control invariant sets $\mathcal{C}_{i}\subseteq\widehat{\mathcal{X}}$, which (by their very construction; see, for example, Section 10.4 in [5]) guarantee the recursive feasibility of our second layer. We stress the fact that (as previously indicated in Section 6) the sets $\mathcal{C}_{i}$ are not computed explicitly, with their existence being a desirable and natural consequence of our proposed approach.
@@ 7.3 Simulation Scenario and Results
To test the obtained control laws, we simulate a scenario in which $\Delta p_{0}[k]=T_{s}v_{0}[k]$, where
@@
is the speed of the virtual $0^{\text{th}}$ car and $\mathds{1}[k]$ denotes the discrete-time Heaviside step function. As can be seen in Figure 3 below, in which $v_{0}[k]$ is depicted in orange, the speed of each car in the platoon will try to match that of the $0^{th}$ car, in order to achieve steady-state evolution.
@@
Moreover, as shown in Figure 4 below, our two-layer control architecture successfully maintains all interdistance variables within their prescribed bounds, thereby avoiding collisions and preventing the cars from drifting too far apart (which can lead to traffic congestion or communication breakdown between consecutive vehicles).
@@
Given the limitations placed upon the signals sent to the car actuators, our control scheme is once again successful. As can be seen from Figures 5 and 6 below, the sum of the signals from the two aforementioned figures is well within the bounds from (50), even taking into account the amplitudes of the entries in $\beta_{s2}[k]$.
@@
The brief intervals in which the two signal vectors are of the same order of magnitude as the first-layer control signals are those in which the second layer intervenes to saturate either the output of the first car’s NRF-based subcontroller (see Figure 8 below), or the speed of the lead car in the platoon (see Figure 9, also located below) with respect to these variables’ prescribed bounds.
@@
Finally, as showcased by the runtime statistics displayed in Tables 3 and 4, which are located at the top of this page, the implementation costs associated with our solution are sufficiently modest such that command computation is well below the sampling period of our control system, namely $T_{s}=0.1$ seconds. Note that all simulations associated with these runtimes were performed in MATLAB R2024b on a mid-range laptop equipped with an Intel Core i9-100885H processor.